# E-values, Multiple Testing and Beyond

### v1

@@ 2.3 E-values and e-BH procedure
A non-negative random variable $E$ is called an e-value if $\mathbb{E}[E]\leq 1$ under the null hypothesis. Suppose we observe $n$ e-values $e_{1},\dots,e_{n}$ corresponding to the hypotheses $H_{1},\dots,H_{n}$. The $\alpha$-level e-BH procedure involves sorting the e-values in decreasing order as $e_{(1)}\geq\dots\geq e_{(n)}$ and rejecting the hypotheses associated with the $\hat{k}$ largest e-values, where $\hat{k}\coloneqq\max\left\{1\leq i\leq n\colon e_{(i)}\geq n/(i\alpha)\right\}$.
Proposition 1 (Theorem 2 of Wang and Ramdas (2022)).
Suppose the e-values satisfy
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq n,$ (3)
where $\mathcal{H}_{0}=\{1\leq i\leq n:\theta_{i}=0\}$. Then, the $\alpha$-level e-BH procedure controls the FDR at the level $\alpha$, regardless of the dependence structure among the e-values.
2.4 Connection between the BH and e-BH procedures
To see the connection between the BH and e-BH procedures, we define the e-value associated with $H_{i}$ to be
$e_{i}=T_{\mathrm{BH}}^{-1}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\},$ (4)
where $T_{\mathrm{BH}}$ is given in (1). By Lemmas 3-4 in Storey et al. (2004), $\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\leq t\}/t$ for $0<t\leq 1$ is martingale with time running backwards with respect to the filtration $\mathcal{F}_{t}=\sigma(\mathbbm{1}\{p_{i}\leq s\}:t\leq s\leq 1,i=1,2,\dots,n)$, and $T_{\mathrm{BH}}$ is a stopping time with respect to $\mathcal{F}_{t}$, where $\mathcal{F}_{t}$ is the sigma field generated by $\mathbbm{1}\{p_{i}\leq s\}$ for $t\leq s\leq 1.$ By the optional stopping time theorem, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]=\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}}{T_{\mathrm{BH}}}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\leq 1\}\right]=n_{0},$
which implies that the e-values defined by equation (4) satisfy (3). Thus by Proposition 1, the corresponding e-BH procedure controls the FDR at the desired level. Moreover, we claim that the e-BH procedure based on the e-values defined in (4) is equivalent to the BH procedure based on $\{p_{1},\dots,p_{n}\}$ in the sense that they produce the same set of rejections.

### v2

@@ 2.2 E-Values and the e-BH Procedure
A non-negative random variable $e$ is called an e-value if it satisfies the condition $\mathbb{E}[e]\leq 1$ under the null hypothesis. Suppose we have $n$ e-values, denoted as $e_{1},e_{2},\dots,e_{n}$, corresponding to the hypotheses $H_{1},H_{2},\dots,H_{n}$. The $\alpha$-level e-BH procedure (Wang and Ramdas, 2022) involves sorting the e-values in decreasing order $e_{(1)}\geq e_{(2)}\geq\dots\geq e_{(n)},$ and rejecting the hypotheses associated with the $\hat{k}$ largest e-values, where $\hat{k}\coloneqq\max\left\{1\leq i\leq n\colon e_{(i)}\geq n/(i\alpha)\right\}$.
Let $\mathcal{H}_{0}=\{1\leq i\leq n\colon\theta_{i}=0\}$ be the set of true null hypotheses. According to Theorem 2 of Wang and Ramdas (2022), a key advantage of the e-BH procedure is that it controls the FDR at level $\alpha$, even when the e-values exhibit arbitrary dependence.
Proposition 1 (Wang and Ramdas (2022), Theorem 2).
Suppose the set of e-values $\{e_{i}\}_{1\leq i\leq n}$ satisfies
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq n.$ (2)
Then, the e-BH procedure controls the FDR at level $\alpha$.
In the context of multiple testing, the requirement that $\mathbb{E}[e]\leq 1$ in the definition of e-values can be relaxed. Specifically, throughout the rest of this paper, we refer to $\{e_{i}\}$ as a set of e-values if they satisfy Condition (2).
2.3 Connection Between Multiple-Testing Procedures and the e-BH Procedure
Given the threshold $T$ defined in (1), the e-BH procedure, defined based on the e-values $e_{i}=\bigl{(}nR_{i}(T)\bigr{)}/m(T)$ for $1\leq i\leq n$, is equivalent to the multiple testing procedures presented in Table A.1 with the same $m(\cdot)$ and $R_{i}(\cdot)$ functions (Li and Zhang, 2025). Here, equivalence means that they produce the same set of rejections.
Unlike p-values, which are defined in terms of tail probabilities, e-values are defined through expectations, making them easier to aggregate or combine. For instance, the arithmetic mean of multiple e-values remains a valid e-value. Building upon the insight that the BH and BC procedures, along with their generalized versions, are equivalent to the e-BH procedure when based on specific forms of e-values, we develop novel multiple testing methods by aggregating e-values derived from different procedures or by assembling e-values obtained from the same procedure applied to distinct subsets of data. By ensuring that these aggregated or assembled e-values satisfy Condition (2), the resulting e-BH procedures maintain finite-sample FDR control. We illustrate these developments with several concrete applications in Sections 3-5.
3 Assembling E-Values Across Data Subsets
In this section, we consider a scenario that we have $L$ sets of e-values, $\{e_{i}^{l}:i\in\mathcal{G}_{l},|\mathcal{G}_{l}|=n_{l}\}$, derived from $L$ distinct datasets, where $\bigcup_{l}\mathcal{G}_{l}=[n]$ and $\mathcal{G}_{l_{1}}\cap\mathcal{G}_{l_{2}}=\emptyset$ for $l_{1}\neq l_{2}$. Each $e_{i}^{l}$ is associated with the hypothesis $H_{i}$, and $\sum_{i\in\mathcal{G}_{l}\cap\mathcal{H}_{0}}\mathbb{E}[e_{i}^{l}]\leq n_{l}$. In this context, our objective is to combine the $L$ sets of e-values into a single e-value vector $(e_{1},\dots,e_{n})$ that satisfies Condition (2).
3.1 Simultaneous Group-Wise and Overall FDR Control
Recall that $\theta_{i}\in\{0,1\}$ represents the true state of hypothesis $H_{i}$, and $\delta_{i}\in\{0,1\}$ denotes the decision rule for $H_{i}$. We define the group-wise FDP and FDR based on $\bm{\delta}$ as follows:
$\mathrm{FDP}_{l}(\bm{\delta})=\frac{\sum_{i\in\mathcal{G}_{l}}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i\in\mathcal{G}_{l}}\delta_{i}},\quad\text{FDR}_{l}(\bm{\delta})=\mathbb{E}[\text{FDP}_{l}(\bm{\delta})],\quad l=1,2,\dots,L.$
A decision rule $\bm{\delta}$ with target FDR level $\alpha$ is said to simultaneously control both group-wise and overall FDR if it uniformly controls the group-wise FDRs for all $1\leq l\leq L$ and maintains the overall FDR at level $\alpha$. Specifically, this means that $\max_{1\leq l\leq L}\text{FDR}_{l}(\bm{\delta})\leq\alpha$ and $\text{FDR}(\bm{\delta})\leq\alpha$.
A decision rule that simultaneously controls both group-wise and overall FDR is relevant to predictive parity within the classification context in the fairness community (Chouldechova, 2017). We compare our definitions with predictive parity in Supplement D.
We propose a multiple testing procedure that simultaneously controls both group-wise and overall FDR by assembling the e-values from the BC procedure applied to each group separately. Specifically, we implement the BC procedure at the level $\alpha$ for each individual group and let
$T_{l}=\sup\left\{0<t<0.5\colon\frac{1+\sum_{i\in\mathcal{G}_{l}}\mathbbm{1}\{p_{i}\geq 1-t\}}{1\vee\sum_{i\in\mathcal{G}_{l}}\mathbbm{1}\{p_{i}\leq t\}}\leq\alpha\right\}$ (3)
be the rejection threshold for the $l$th group with $1\leq l\leq L$. Define
$e_{i}=\frac{n_{l}w_{i}\mathbbm{1}\{p_{i}\leq T_{l}\}}{1+\sum_{j\in\mathcal{G}_{l}}\mathbbm{1}\{p_{j}\geq 1-T_{l}\}},$ (4)
for $i\in\mathcal{G}_{l}$, where $w_{i}>0$ represents the weight for the $i$th hypothesis, which will be specified in Section 3.1.1. After collecting the e-values from each group, we implement the e-BH procedure at level $\alpha$. The testing procedure is summarized in Algorithm 1.
Algorithm 1 Multiple testing procedure that simultaneously controls both group-wise and overall FDR
1:p-values $p_{1},\dots,p_{n}$, group indices $\mathcal{G}_{1},\dots,\mathcal{G}_{L}$, significance level $\alpha$
2:for $l=1,\dots,L$ do
3:  Implement the BC procedure utilizing the p-values $\{p_{i}\colon i\in\mathcal{G}_{l}\}$ at the level $\alpha$.
4:  Calculate the threshold $T_{l}$ using (3).
5:  for $i\in\mathcal{G}_{l}$ do
6:   Calculate the e-value $e_{i}$ using (4).
7:  end for
8:end for
9:Assemble the e-values from all groups.
10:Run the e-BH procedure utilizing the assembled e-values at the level $\alpha$.
11:The indices of rejected hypotheses.
3.1.1 Choosing Weights and FDR Control
Controlling the overall FDR requires that the e-values defined in (4) satisfy Condition (2). One approach is to set $w_{i}=1$ for all $i$. Alternatively, the group size can be taken into account by setting $w_{i}=n/(Ln_{l})$ for all $i\in\mathcal{G}_{l}$ and $l=1,\dots,L$. According to Proposition 6 in Li and Zhang (2025), both strategies satisfy Condition (2). The e-BH procedures based on these weight choices are referred to as eBH_1 and eBH_2, respectively. However, our simulations indicate that eBH_1 and eBH_2 often suffer from low statistical power. To enhance efficiency, we propose using a data-dependent weight approach inspired by the leave-one-out technique (Barber et al., 2020).
Denote the p-values in the $l$th group by $\mathbf{p}_{l}=\{p_{i}\}_{i\in\mathcal{G}_{l}}$. Write $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$, and let $\mathbf{p}_{l,i}$ for $i\in\mathcal{G}_{l}$ be the collection of p-values obtained by replacing $p_{i}$ with $\tilde{p}_{i}$ in $\mathbf{p}_{l}$. By viewing $T_{l}$ as a functions of $\mathbf{p}_{l}$, we define $T_{l,i}=T_{l}(\mathbf{p}_{l,i})$, i.e., the threshold of the BC procedure applied to the set of p-values $\mathbf{p}_{l,i}$. We define the data-dependent weights as
$w_{i}=\frac{\frac{n}{n_{l}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{l}}\mathbbm{1}\{p_{j}\geq 1-T_{l}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{l}}\mathbbm{1}\{p_{j}\geq 1-T_{l}\}\right)+\sum_{l^{\prime}\neq l}\sum_{j\in\mathcal{G}_{l^{\prime}}}\mathbbm{1}\{p_{j}\geq 1-T_{l^{\prime},j}\}},$ (5)
for $i\in\mathcal{G}_{l}$. The e-BH procedure, based on the weights specified in (5), will henceforth be referred to as the eBH_Ada method in the following discussions. If the null $p$-values satisfy the following condition:
$P(p_{i}\leq a)\leq P(p_{i}\geq 1-a)=P(1-p_{i}\leq a),\quad\text{ for all }0\leq a\leq 0.5,$ (6)
then eBH_Ada has finite-sample FDR control.